# Identity-related Speech Suppression in Generative AI Content Moderation

### v1

@@ 3.2 Speech suppression measures for numerical scores
4 Existing and New Datasets
Dataset Data subset Total Avg. Len.
items (words)
Jigsaw Kaggle - 445293 61
Jigsaw Bias - 60560 5
Stormfront - 10944 18
TweetEval Hate 12962 21
Offensive 14100 22
OpenAI - 1680 111
Movie Plots - 6001 612
TV Synopses Short TMDB 14760 44
Medium Wiki 6603 94
Long IMDB 1843 867
Traditional - 543859 52
GenAI - 30887 218
Table 1: Datasets included in the benchmark, with the total number of items from the dataset we include in the benchmark and the average word length per item.
@@ 4 Existing and New Datasets
We assess the automated content moderation APIs and build a benchmark from seven datasets; five from previous work and two introduced in this work. [^1] All datasets, including the added identity categorization and the associated code, are available at: https://github.com/genAIaudits/speech-suppression. The five datasets from previous work were chosen because of their previous use assessing content moderation systems (see, e.g., [22]). Most datasets we include in this benchmark include Boolean classification labels indicating whether the provided text snippet should be categorized as hateful, toxic, or otherwise offensive. These datasets were created to help automatically monitor online comment sections and other short-form content (e.g., tweets), and are generally reflective of these goals. These are the Jigsaw Kaggle [6, 5], Jigsaw Bias [10], Stormfront [8], and TweetEval [2] datasets. The Jigsaw Bias and Stormfront dataset labels indicate whether the text is toxic, the TweetEval dataset provides a subset labeled to indicate hate and another labeled for offensive content, and the Jigsaw Kaggle dataset provides content categorized based on an overall toxicity flag and six harmful content types (severe toxic, obscene, threat, insult, identity hate, and sexually explicit). All these datasets contain text items that are fairly short, with an average length between 5 and 61 words. Jigsaw Kaggle’s dataset includes about $2$ million text instances; we include only the subset of about $445,000$ that were also manually identified for association with an identity group. See Table 1 for dataset summary statistics; when considered as a group, we identify all these datasets (Jigsaw Kaggle, Jigsaw Bias, Stormfront, and TweetEval) as “Traditional” (i.e., pre-generative AI content moderation datasets).
@@
Dataset Data subset Non-wh. White Men Women Christ. Non-chr. LGBT Str. Disab.
@@
Stormfront - 476 255 1025 727 118 327 91 1 18
TweetEval Hate 932 135 2611 4805 121 332 90 1 27
Offensive 303 58 3321 2572 101 84 73 2 56
OpenAI - 184 46 567 482 38 101 117 5 32
Movie Plots - 1073 269 351 2675 245 211 394 412 470
TV Synopses Short TMDB 45 20 20 142 416 158 613 425 168
Medium Wiki 60 10 20 82 216 63 338 240 92
Long IMDB 47 11 30 143 83 23 223 128 47
Traditional - 43838 30738 57100 67892 49367 39645 27457 4460 10165
GenAI - 1409 356 988 3524 998 556 1685 1210 809
@@
Table 2: The number of instances per dataset and identity group after identity categorization is performed.
More recent datasets have been developed to test generative AI content moderation systems, including OpenAI’s content moderation dataset [22] that includes longer-form text tagged with information about whether it should be categorized as any of nine types of violating content (sexual, hate, violence, harassment, self-harm, self-harm with intent, sexual content relating to minors, threatening hate speech, or graphic violence). [21] also develop a dataset of longer-form text containing television episode synopses (of varying lengths), however the dataset is fairly small (1,392 episodes). Inspired by their focus on cultural content, we introduce additional datasets containing television and movie plots.
4.1 New datasets: Television and movie plots
In order to test longer-form creative content of the type relevant to generative text systems and our associated concerns about speech suppression, we developed datasets of television episode and movie plot synopses. Using The Movie Database (TMDB [^2] https://www.themoviedb.org/), we gathered the top $10,000$ television shows and movies as of Summer 2024 and filtered the lists to only include English-language shows and movies released in the United States. For each television show, overviews for all episodes in the first season were collected. The show and movie names were then identified on Wikipedia, where episode synopses and movie plots were collected. Longer user-generated summaries of TV episodes were additionally collected from IMDB. The summary statistics for the OpenAI, Movie Plots, and TV synopses data, collectively referred to as the GenAI datasets, are also given in Table 1.
We also collected age ratings from TMDB for movies (G, PG, PG-13, R, and NC-17) and from IMDB for television episodes (TV-Y, TV-Y7, TV-G, TV-PG, TV-14, TV-MA). These age ratings are established by an external organization (the Motion Picture Association) and indicate the maturity level of the content; for example, PG-13 content is considered appropriate for people aged 13 and up. Using these age ratings, we constructed two sets of labels for each TV episode or movie: PG ok and PG-13 ok. The PG ok labels classify an episode or movie as appropriate, or not in need of content moderation, if it’s rated G or PG (respectively for TV shows, TV-Y, TV-Y7, TV-G, or TV-PG), and the PG-13 ok labels mark G, PG, or PG-13 (for TV shows, TV-Y, TV-Y7, TV-G, TV-PG, or TV-14) rated episodes or movies as appropriate. Unrated episodes or movies were excluded from the final dataset. We claim these labels are best interpreted in the context of a false positive score or other measures that focus on incorrectly flagging content as violating; since we include only episode and movie synopses in our data, and not, e.g., the full scripts, there is likely to be content missing from the synopses that is included in the full episode or movie as rated. Under a conservative interpretation of these ratings for our PG-13 ok labels, the movie synopses for PG-13 rated movies are certainly appropriate (i.e., should not be flagged), and higher rated movies may also have appropriate synopses since TMDB, Wikipedia, and IMDB are unlikely to contain inappropriate text even for shows and movies with higher age ratings.
4.2 Identity categorization
To differentiate content moderation success rates by identity group, we tag each instance of text in each dataset with identity groups, where each item could be tagged with multiple identity groups or none. Text is associated with specific identities (e.g., Black or lesbian) and then grouped with related identities into general identity groups (e.g., non-white or LGBT) in order to create groups large enough to assess content moderation trends. Most datasets do not already include identity attributes, so we create these associations following two main strategies: text-based identification of explicit references to an identity group, and external association of cultural content with an identity group. Additionally, the Jigsaw Bias dataset is created by substituting identity terms into various repeated phrases (e.g., “hug gay”, “hug male”, and so on) so those identity groups are directly extracted, and the Jigsaw Kaggle dataset includes manually identified labels [5, 6] for a subset of about 445K instances which we use directly.
@@ 4.2 Identity categorization
In order to identify explicit references to an identity group in text, we create lists of identity-related terms that include both slurs or slang (Appendix Tables 5 and 6) and neutral descriptors (Appendix Table 7) about an identity group. These lists were collected from Wikipedia (e.g., Wikipedia’s “List of ethnic slurs”), and then curated to remove terms that are also commonly used words in other contexts (e.g., “black”) so as to conservatively identity-tag text that is actually likely referring to a given identity group. Text from the Stormfront, TweetEval, and OpenAI datasets were then categorized as associated with a specific identity if they included any of these identity-related terms.
@@
We were able to validate this identity categorization scheme using the Jigsaw Kaggle data’s manually identified labels, and found that the auto-tagging scheme had high accuracy when identifying text about the Christian (96%), non-Christian (98%), white (94%), non-white (94%), straight (99.8%), LGBT (99%), disability (98%), and women (92%) identity categories, and had lower accuracy (77%) in identifying text about men, largely due to false negatives (i.e., missing tags for text that should have been identified as about men). The false positive rates (i.e., text incorrectly tagged as relating to an identity group), were all below 4%, except for women (9%) and men (25%).
@@
The movie and television datasets were associated with identity groups based on external information about the show or movie. Wikipedia has categories that pages can be tagged with (e.g., “Category:LGBT-related films”), and these were used to build up a list of movies and television shows associated with specific identities. The Wikipedia categories and other URLs used to create identity-related lists of movies and TV shows can be found in Appendix Tables 8 and 9. In order to identify movies and television shows associated with some dominant groups where such categories did not exist (e.g., television shows about white people), larger sets of shows were collected (e.g., television shows set in Europe) and then shows identified through a different category list as not belonging to the dominant group (e.g., shows about non-white people) were removed. Additionally, user-generated tags associated with television episodes on IMDB were used to add further identity tags to episodes, using the procedures and tag lists of [21].
@@
Counts of the number of instances associated with each general identity group per dataset resulting from this process are given in Table 2. While identity groups have a low number of resulting tagged instances for some datasets—for example, very few text instances are tagged as “straight” following the text reference identification method used for the Stormfront, TweetEval, and OpenAI datasets—when aggregated for the traditional and generative AI datasets, each identity group makes up at least approximately 1% of each dataset. The group with the lowest identified representation in the traditional dataset is straight people (0.8%) and in the generative AI dataset is white people (1%), while the highest represented group in both datasets is women (12% traditional, 11% generative AI).
5 Experiments
We are interested in assessing content moderation APIs for identity-related speech suppression and specifically in answering the following experimental research questions:
  1. 1.
  2. 2.
  3. 3.
    Do APIs have different behaviors on creative generative AI text than on traditional user-generated short-form text?
We first describe the five publicly available content moderation APIs that we assess in this work. Code to replicate these experiments can be found at: https://github.com/genAIaudits/speech-suppression.
5.1 Automated content moderation APIs
API name & version Categorization Aims Return Value(s)
Perspective v1alpha1 toxicity score
Anthropic claude-3 -haiku- 20240307 violence, illegal activities, hate speech, explicit sexual content, harmful misinfo./conspiracy theories overall flag
OpenAI Moderation Endpoint 007 sexual, hate, violence, harassment, self-harm, self-harm/intent, sexual/minors, hate/threatening, violence/graphic per-cat. scores, per-cat. flags, overall flag
Llama Guard 2-8b via OctoAI violence and hate, sexual content, criminal planning, guns/illegal weapons, self-harm, regulated/contr. substances per-category flags, overall flag
Google Moderate text July 2024 toxic, insult, profanity, derogatory, sexual, death and harm/tragedy, violent, firearms/ weapons, public safety, health, religion and belief, illicit drugs, war and conflict, politics, finance, legal per-category scores
Table 3: The automated content moderation APIs and associated characteristics.

### v3

@@ 3.2. Speech suppression measures for numerical scores
4. Audit methodology
In order to assess the identity-related speech suppression due to automated content moderation systems across differing types of content, identities, and APIs, we develop an audit methodology based on three key steps:
  1. (1)
    content moderation texts from both traditional online moderation and generative AI contexts, with labels indicating if content should be allowed;
  2. (2)
    identity categorization of each text as associated with each of nine identities: men, women, christian, non-christian, straight, LGBT, white, non-white, or disability; and
  3. (3)
    running each of five publicly available content moderation APIs.
Figure 1. Overview of the audit methodology pipeline. Dataset information shown includes total number of instances ($n$) and average number of words per instance. Content moderation API information shown includes the version number audited, return types, and categorization aims.
[…]
[Overview of the audit methodology pipeline]Audit methodology pipeline parts are: 1. identity tagging, including template-based, individually coded, text references, and external associations as sub-methods; 2. datasets, including traditional content moderation datasets and generative AI datasets; 3. content moderation APIs including traditional content moderation context APIs and those from a generative AI context, and; 4. outputs, including both scores and boolean flags. Arrows indicate flow from one part of the pipeline to the next.
An overview of the audit methodology can be found in Figure 1 and data and code to replicate these experiments can be found at: https://github.com/genAIaudits/speech-suppression.
4.1. Datasets
We build an automated content moderation benchmark from seven datasets; five from previous work and two introduced in this work. The five datasets from previous work were chosen because of their previous use assessing content moderation systems (see, e.g., (markov2023holistic, )). These previous datasets include Boolean classification labels indicating whether the provided text snippet should be categorized as hateful, toxic, or otherwise offensive as determined by human raters. As we will discuss below, we introduce a new methodology for creating such labels in the context of speech suppression that takes advantage of its focus on incorrect flagging behavior, and use this methodology to build on existing human-curated information to label the two new data sets.
Traditional datasets
@@ 4.1. Datasets
Four of the included datasets were created to help automatically monitor online comment sections and other short-form content (e.g., tweets), and are generally reflective of these goals. These are the Jigsaw Kaggle (jigsaw-kaggle-unintended-bias, ; borkan2019nuanced, ), Jigsaw Bias (dixon2018measuring, ), Stormfront (de2018hate, ), and TweetEval (barbieri2020tweeteval, ) datasets. The Jigsaw Bias and Stormfront dataset labels indicate whether the text is toxic, the TweetEval dataset provides a subset labeled to indicate hate and another labeled for offensive content, and the Jigsaw Kaggle dataset provides content categorized based on an overall toxicity flag and six harmful content types (severe toxic, obscene, threat, insult, identity hate, and sexually explicit). All these datasets contain text items that are fairly short, with an average length between 5 and 61 words. Jigsaw Kaggle’s dataset includes about $2$ million text instances; we include only the subset of $445,293$ that were also manually identified for association with an identity group. See Figure 1 for dataset summary statistics; when considered as a group, we identify all these datasets (Jigsaw Kaggle, Jigsaw Bias, Stormfront, and TweetEval) as “Traditional” content moderation datasets.
Generative AI data
More recent datasets have been developed to test generative AI content moderation systems. This includes OpenAI’s content moderation dataset (markov2023holistic, ) we include in our benchmark that contains longer-form text tagged with information about whether it should be categorized as any of nine types of violating content (sexual, hate, violence, harassment, self-harm, self-harm with intent, sexual content relating to minors, threatening hate speech, or graphic violence). (mahomed2024auditing, ) also develop a dataset of television episode synopses, however the dataset is fairly small (1,392 episodes); inspired by their focus on cultural content, we introduce additional datasets containing television and movie plots.
4.1.1. New datasets: Television and movie plots
In order to test longer-form creative content of the type relevant to generative text systems and our associated concerns about speech suppression, we developed datasets of television episode and movie plot synopses. Using The Movie Database (TMDB)(tmdb, ), we gathered the top $10,000$ television shows and movies as of Summer 2024 and filtered the lists to only include English-language shows and movies released in the United States. For each television show, overviews for all episodes in the first season were collected. The show and movie names were then identified on Wikipedia, where episode synopses and movie plots were collected. Longer user-generated summaries of TV episodes were additionally collected from IMDB. The summary statistics for the OpenAI, Movie Plots, and TV synopses data, collectively referred to as the GenAI datasets, are also given in Figure 1.
It can be difficult for researchers outside of an industry context to create new datasets for content moderation tasks, since many of the previously developed traditional datasets were hand-labeled by contract or ”gig” workers and this may be cost prohibitive for the development large datasets in an academic context. Our focus on speech suppression allows a wider range of collected data to be useful to our task; as we describe further in what follows, we were able to automatically generate labels based on television or movie age ratings, taking advantage of previous human labor applicable to our goals.
We collected age ratings from TMDB for movies (G, PG, PG-13, R, and NC-17) and from IMDB for television episodes (TV-Y, TV-Y7, TV-G, TV-PG, TV-14, TV-MA). These age ratings are established by an external organization (the Motion Picture Association) and indicate the maturity level of the content; for example, PG-13 content is considered appropriate for people aged 13 and up. Using these age ratings, we constructed two sets of labels for each TV episode or movie: PG-ok and PG-13-ok. The PG-ok labels classify an episode or movie as appropriate, or not in need of content moderation, if it’s rated G or PG (respectively for TV shows, TV-Y, TV-Y7, TV-G, or TV-PG), and the PG-13-ok labels mark G, PG, or PG-13 (for TV shows, TV-Y, TV-Y7, TV-G, TV-PG, or TV-14) rated episodes or movies as appropriate. Unrated episodes or movies were excluded from the final dataset. Additionally, since PG-13 ratings were not introduced for movies until July, 1984 and age ratings were not introduced for television shows until January, 1997, we include only movies released after 1985 and television shows released after 1997. We will report detailed results on both PG-ok and PG-13-ok labels, and where aggregating across datasets will use the PG-13-ok labels for these datasets.
These introduced age-based labels align well with our speech suppression measurement goals, since they focus on incorrectly flagging content as violating, and would be less appropriate to use to identify content that should be flagged. Since we include only episode and movie synopses in our data, and not, e.g., the full scripts, there is likely to be content missing from the synopses that is included in the full episode or movie as rated. Under a conservative interpretation of these ratings for our PG-13-ok labels, the movie synopses for PG-13 rated movies are certainly appropriate for kids 13 and older (i.e., should not be flagged), and higher rated movies may also have appropriate synopses since TMDB, Wikipedia, and IMDB are unlikely to contain inappropriate text even for shows and movies with higher age ratings. We note that using the PG-13-ok labels makes sense for all AI services, since they include a requirement in their terms of use that users must be at least 13 years old (openai2024termsofuse, ), or in some cases at least 18 years old (anthropic2024termsofuse, ), likely to comply with COPPA, a U.S. law that imposes additional restrictions on websites directed to children under 13 (coppa, ).
4.2. Identity categorization
@@
Dataset Data subset Non-white White Men Women Christian Non-Christian LGBT Straight Disability
@@
Stormfront - 476 255 648 536 118 327 91 1 18
TweetEval Hate 932 135 1956 4535 121 332 90 1 27
TweetEval Offensive 303 58 779 780 101 84 73 2 56
OpenAI - 184 46 384 400 38 101 117 5 32
Movie Plots - 1000 239 330 2417 213 183 351 383 439
TV Synopses Short TMDB 45 20 20 110 338 115 584 332 114
TV Synopses Medium Wiki 60 10 20 50 188 53 324 189 74
TV Synopses Long IMDB 26 11 30 137 70 17 207 94 34
Traditional - 43838 30738 53526 65639 49367 39645 27457 4460 10165
GenAI - 1315 326 784 3114 847 469 1583 1003 693
@@
Table 1. The number of instances per dataset and identity group after identity categorization is performed.
In support of our goal to assess identity-related bias in speech suppression, we need each text to be associated with zero, one, or multiple identity groups. Specifically, we aimed to determine if each text instance should be associated with identity categories based on race, gender, religion, sexual orientation, and disability. The choice of these identity categories is motivated by U.S. non-discrimination law (they are protected categories under civil rights laws and interpretations) and are common choices in the literature, including the literature and datasets we directly build upon (see, e.g., (borkan2019nuanced, ; dixon2018measuring, ; mahomed2024auditing, )). In what follows, we will group identities to compare dominant versus non-dominant groups, as is common practice in quantitative work for handling marginalized groups’ small sample sizes, and qualitative work comparing dominant group outcomes to others (see e.g., (gillespie2024generative, ; 10.1145/3449131, ; 10.1145/3449100, )).
Only two of the datasets in our benchmark were previously tagged with identity information; the Jigsaw Bias and Jigsaw Kaggle datasets. The Jigsaw Bias dataset (dixon2018measuring, ) is template-based, substituting identity terms into various repeated phrases (e.g., “hug gay”, “hug male”, and so on) so we can directly extract those identity groups. The Jigsaw Kaggle dataset includes human individually coded identity labels (borkan2019nuanced, ; jigsaw-kaggle-unintended-bias, ) for a subset of about 445K instances; we directly use these individually coded identities.
The remaining five datasets do not already include identity attributes, so we create these associations following two main strategies: identification of explicit text references to an identity group, and external association of cultural content with an identity group. For both introduced methodologies, we follow the strategy of associating text with specific identities (e.g., Black or lesbian) and then grouping those identities into general identity groups (e.g., non-white or LGBT) in order to create groups large enough to assess content moderation trends. This strategy of going from specific identities to larger more general identity groups allows us to categorize text according to fine-grained information (e.g., a slur specifically used about lesbians) while still considering trends over a larger sized group.
4.2.1. Text references
In order to automatically tag text as related to an identity group, we determined that we would need a wide variety of terms describing identity groups to draw from. Once we have an appropriate list of such terms, we can then tag any text that explicitly references a group as related to that identity. We turned to Wikipedia for pages identifying slurs and slang terms associated with various identity groups. We believe that Wikipedia’s wide range of editors from a variety of backgrounds and its consensus-building joint editorial process make it useful for building out identity-related information, such as these lists of slurs or slang terms. Related work has found that Wikipedia content on taboo topics is of relatively higher quality, and that community-driven Wikipedia curation can be useful in the context of AI evaluation (champion2023taboo, ; kuo2024wikibench, ).
@@ 4.2.1. Text references
We created lists of identity-related terms that include both slurs or slang (Appx. Tables 6 and 7) and neutral descriptors (Appx. Table 8) about an identity group. These lists were collected from Wikipedia (e.g., Wikipedia’s “List of ethnic slurs”), and then curated to remove terms that are also commonly used words in other contexts (e.g., “black”) so as to conservatively identity-tag text as referring to a given identity group. Text from the Stormfront, TweetEval, and OpenAI datasets were then categorized as associated with a specific identity if they included any of these identity-related terms.
Validation
@@
We were able to validate this identity categorization scheme using the Jigsaw Kaggle data’s manually identified labels, and found that the auto-tagging scheme had high accuracy when identifying text across all identity categories, including Christian (96%), non-Christian (98%), white (94%), non-white (94%), straight (99.8%), LGBT (99%), disability (98%), women (95%), and men (93%). Interestingly, our validation confirmed that by building from the identified Wikipedia entries, and despite this seemingly simplistic identity-tagging scheme, we were able to closely match human judgment in identifying text as related to a variety of identities.
4.2.2. External associations
@@
The movie and TV datasets were associated with identity groups based on external information about the show or movie. Similarly to the use of Wikipedia to identify slurs and slang, we found Wikipedia and IMDB’s crowd-sourced lists and tags related to identity and media to be usefully varied across identity and detailed in composition. Wikipedia has categories that pages can be tagged with (e.g., “Category:LGBT-related films”), and these were used to build up a list of movies and shows associated with specific identities (specific categories and URLs are given in Appx. Tables 9 and 10). In order to identify movies and shows associated with some dominant groups where such categories did not exist (e.g., TV shows about white people), larger sets of shows were collected (e.g., TV shows set in Europe) and then shows identified as not belonging to the dominant group (e.g., shows about non-white people) were removed. Additionally, user-generated tags associated with television episodes on IMDB were used to add further identity tags to episodes, using the procedures and tag lists of (mahomed2024auditing, ).
We note here that we marked TV and movies as related to an identity based on salience as indicated by crowd-sourcing projects as opposed to, e.g., based on the percent of actors portraying or from a specific identity. This is a purposeful choice based on an understanding from media studies of the important difference between genuine representation and casting choices. As one example, this distinction has been codified as the “Bechdel Test” wherein a movie is considered to have good representation of women if it has two named women characters who talk to each other about something other than a man (bechdeltest, ); we use a crowd-sourced site with a listing of movies that pass the Bechdel Test to identify movies relating to women (see Appx. Table 9). More broadly, we claim that movies and television shows individually identified by Wikipedia editors or IMDB users as related to an identity are likely to be genuinely salient for that group. [^1] In fact, instructions for Wikipedia categories make some of this explicit. For example, the Wikipedia category of “African-American Films” (https://en.wikipedia.org/wiki/Category:African-American_films) explicitly states that “This category is for Films identified by reliable sources as “African-American films”. This should not be used as a catch-all for all films starring or made by African-Americans. Nor should it be used as an umbrella category for all films about African Americans.”
4.2.3. Identity-tagging results
@@
Based on all four of these identity-related tagging methodologies, counts of the number of instances associated with each general identity group per dataset resulting from this process are given in Table 1. While identity groups have a low number of resulting tagged instances for some datasets—for example, very few text instances are tagged as “straight” following the text reference identification method used for the Stormfront, TweetEval, and OpenAI datasets—when aggregated for the traditional and generative AI datasets, each identity group makes up at least approximately 1% of each dataset. The group with the lowest identified representation in the traditional dataset is straight people (0.8%) and in the generative AI dataset is white people (1.4%), while the highest represented group in both datasets is women (12% traditional, 13% generative AI). Numerically, given the large size of the datasets assembled, the representation even of the smallest group is still fairly large, with $4,460$ text instances associated with straight identity marked in the traditional dataset and $326$ instances associated with white identity in the generative AI dataset.
4.3. Automated content moderation APIs
We identified five publicly available automated content moderation APIs to audit covering a variety of types and goals. An overview of these systems, including the specific version number audited, can be found in Figure 1. Jigsaw’s Perspective API is a traditional content moderation system used for moderation of online comment sections and other user-generated content. Llama Guard (accessed via Octo AI’s API platform (octoaiAPI, )) and Anthropic (anthropicAPI, ) provide specific prompts (see Appx. Figures 4, 5, and 6) to give to generative AI systems. In other words, these are large language models being used for a content moderation purpose. Anthropic states the aims of this system as focused on user-generated text (i.e., a traditional content moderation goal), and the language model used is their general model (Claude) (anthropicAPI, ). Llama Guard, in contrast, focuses on generative AI conversation use cases and has fine-tuned Llama for the purpose of content moderation in this generative AI context (inan2023llama, ). OpenAI (markov2023holistic, ) and Google (googleAPI, ) have automated content moderation APIs that are separately created models aimed to be used to filter content. Google’s moderation API documentation refers generally to the goal of moderating content, regardless of the source (i.e. whether AI- or user-generated). OpenAI’s moderation endpoint (markov2023holistic, ) is used on the generated outputs of their AI systems (e.g., GPT-4) such that what is visible to a user of their standard web interface is the result of running the AI API followed by the moderation endpoint. This is our main use case of concern in a generative AI context; the speech that will be blocked before being shown to a user.
Each system is designed for slightly different categorization aims, from toxicity to sexual content to politics (see Fig. 1). Most systems identify a broad set of categorization aims; the only exception is Jigsaw’s Perspective which is solely focused on identifying toxicity. For the remaining APIs, a few categories are consistently considered: toxicity or hate, violence, sexual content, and self-harm. Llama Guard, Anthropic, and Google additionally include categories focused on weapons, illegal drugs, and other illegal activities. Beyond these, Google has additional categories that it flags text based on that are not common to the other APIs, with scores provided on a per-category basis for the additional categories of public safety, health, religion and belief, war and conflict, politics, finance, and legal content.
Throughout, we are interested in assessing each APIs behavior from the perspective of incorrect flagging based on the dataset labels. Thus, it is also important to consider the goals of labeling for each of the datasets. All the traditional datasets were labeled with the goal of identifying toxic or hateful speech for the moderation of online user-generated comments. We note that in some cases this includes identifying violence or sexual content as toxic. The OpenAI dataset was labeled according to the OpenAI API moderation aims, i.e., a text instance is labeled as “should flag” if it contains sexual content, hate, violence, harassment, or self-harm (including subcategories such as sexual content involving minors). Finally, recall that the movie and TV datasets were labeled based on age ratings, which are themselves based on identification of violence, sexual content, inappropriate language (e.g., hateful language), and other adult themes (e.g., self-harm and illegal drug use) (mpa2020rules, ). Thus, despite some differences in API aims, we find that the labeling and categorization aims across datasets and APIs are largely focused on identifying toxicity or hate, sexual content, and violence, and provide a reasonably consistent basis for comparison across APIs and datasets.
The types of values returned by these moderation systems also vary, including overall flags indicating content deemed violating, per-category flags indicating content identified with that category, as well as overall or per-category scores (see Fig. 1). Some systems return more than one of these types of values. While Google’s returned per-category scores are meant to represent confidence that an input belongs to the associated category (googleAPI, ), and thus are comparable across categories, OpenAI’s per-category scores must be normalized to be comparable across categories (mahomed2024auditing, ). We follow the normalization mechanism described in (mahomed2024auditing, ); we determine score flagging thresholds (see Appendix Table 11) per category and divide by these thresholds so that scores above $1.0$ flag. While a previous audit of OpenAI determined that multiple runs per instance were useful for the stability of the results (mahomed2024auditing, ), we run each instance through each API only once given the size of the combined datasets and the cost of some APIs.
All instances from each dataset were sent through the five automated content moderation APIs to receive categorization and/or scored results. The resulting benchmark dataset contains $543,859$ instances of traditional content moderation text and $23,722$ generative AI instances, each annotated with zero, one, or more of nine identity groups, a label indicating whether the text should be flagged as violating, and the flag and/or score results from all five content moderation APIs.
5. Findings
Using the audit methodology described in Section 4, we collected API results across benchmark datasets. Our key findings are that:
  1. (1)
    all content moderation systems tested are more likely to incorrectly suppress identity-related text, in many cases at two to three times the rate for non-identity-related text;
  2. (2)
    generative AI datasets have different speech suppression patterns than traditional datasets and these are not accounted for simply by the length of the texts;
  3. (3)
    speech that is incorrectly suppressed on traditional datasets is often political, while generative AI suppression is more often TV violence; and
  4. (4)
    identity-related speech suppression is related to stereotypes or harms particular to an identity group, with, e.g., disability-related content more likely to be flagged for self-harm while non-Christian content flags as hateful.
In what follows we further detail these findings.
5.1. All APIs suffer from identity-related speech suppression